OIC - Applying Concepts Ch 2 #5

5. The patient has two insurance policies. Each policy's coinsurance is 80-20. If the patient has charges totaling $845, what amount should the medical insurance specialist expect to be reimbursed from Insurer A? From Insurer B? For what amount should the patient be billed?
Answer #1

The patient has two insurance policies, and both policies' coinsurance is 80-20. So, it does not matter which Insurer will reimburse first.

Let's assume; The patient will use his insurance A first, then insurance B.

Patient's total bill = $845.

Insurer A will pay = (80/100)*845 = $676.

So, the left amount of the bill is = (845-676) = $169.

Now, The Insurer B is going to pay the 80% of $169 that means = (80/100)*169 = $135.2.

So, the amount will be billed for patient is = (169-135.2) = $33.8.


answered by: Ctiverda